# Input Convex Lipschitz RNN: A Fast and Robust Approach for Engineering Tasks

### v3

@@ III-B2 LMPC formulation
where $S(\triangle)$ denotes the set of piecewise constant functions with period $\triangle$, $\tilde{x}$ denotes the predicted state trajectory, and $N$ denotes the number of sampling periods in the prediction horizon. The objective function $\mathcal{L}$ outlined in Eq. (3a) integrates a cost function $J$ that depends on both the control actions $u$ and the system states $x$. The system dynamic function $F_{nn}(\tilde{x}(t),u(t))$ expressed in Eq. (3b) is parameterized by an RNN (e.g., the proposed ICLRNN in this work). Eq. (3c) encapsulates the constraint function $U$, delineating feasible control actions. Eq. (3d) establishes the initial condition $\tilde{x}(t_{k})$ in Eq. (3b), referring to the state measurement at $t=t_{k}$. Lastly, Eq. (3e) represents the Lyapunov-based constraint $V$, ensuring closed-loop stability within LMPC by mandating a decrease in the value of $V(x)$ over time.
Remark 1 ([9]).
A convex neural network-based LMPC remains convex even when making multi-step ahead predictions, where the prediction horizon is greater than 1. This statement holds true if we incorporate an inherently input convex neural network (e.g., ICLRNN) to the LMPC, as described in Eq. (3b), and if we guarantee that the cost function $J$ in Eq. (3a) is input convex throughout the task.
IV Empirical Evaluation
@@
Moreover, since linear systems have been well studied with numerous classic data-driven modeling approaches available, making the utilization of neural networks unnecessary in general. However, learning nonlinear dynamics presents a different challenge, and many traditional data-driven modeling techniques fail to obtain an accurate nonlinear dynamic model. Therefore, we choose the Continuous Stirred Tank Reactor (CSTR), which is a widely adopted benchmark example in chemical process modeling, due to its high nonlinearity, especially for states significantly distant from the steady-state.
@@
Drawing from the aforementioned considerations, we meticulously choose two case studies to showcase the computational efficiency and non-adversarial robustness of ICLRNN. These include predicting solar irradiance for Solar PV systems using real-time data from LHT Holdings plants, and implementing an NN-MPC for a chemical reactor. It is worth noting that ablation studies are conducted to ascertain the optimal hyperparameter values for all subsequent tasks.
Moreover, given that our research encompasses both Lipschitz-constrained neural networks and Input Convex neural networks, we meticulously select the most suitable approach from these categories and ensure consistency by employing a unified set of baseline methods across both tasks. Our primary aim, however, is to showcase the superiority of our proposed method over any alternatives within these two distinct families.
IV-A Real-World Solar Irradiance Prediction
In this task, we aim to forecast solar irradiance using real-time data sourced from LHT Holdings, a notable wood pallet manufacturing industry based in Singapore (for a detailed manufacturing pipeline of LHT Holdings, refer to Fig. 5 in Appendix B). Before the installation of the Solar photovoltaic (PV) system, LHT Holdings relied solely on the main utility grid to fulfill its energy needs. The Solar PV system was successfully installed by 10 Degree Solar in late 2022, as depicted in Fig. 7 in Appendix B. This transition to solar energy was primarily driven by the aim to minimize manufacturing costs. Three significant uncertainties were identified within the industry, namely solar irradiance affecting the Solar PV system’s efficiency, the dynamic real-time pricing of energy sourced from the main utility grid, and the fluctuating energy demand stemming from unexpected customer orders.
At present, the Solar PV system serves as the primary energy source for the industrial facility, while the main utility grid acts as a secondary energy source to supplement any deficiencies in solar energy production. Any surplus solar energy generated beyond the current requirements is efficiently stored in batteries for future utilization (i.e., see Fig. 6 in Appendix B). Hence, in order to optimize real-time energy planning and mitigate manufacturing costs associated with energy procurement from the main utility grid, a high level of precision and accuracy in forecasting solar irradiance becomes imperative. Moreover, a simple and robust trained NN with relatively high accuracy can be seamlessly integrated into applications like energy dispatch optimization. To achieve these goals, the Solar Energy Research Institute of Singapore (SERIS) and Singapore Institute of Manufacturing Technology (SIMTech) have played pivotal roles by installing various sensors specifically designed for the Solar PV system. These sensors include irradiance sensors, humidity sensors, wind sensors, and module temperature sensors. Drawing on parameters akin to those employed by [30], we utilize minute-based average global solar irradiance, ambient humidity, module temperature, wind speed, and wind direction to predict solar irradiance in the subsequent minute.
In this time-series forecasting task, all models undergo training and evaluation using identical settings. These settings encompass the mean squared error as the loss function, the Adam Optimizer, the ReLU activation function except for LRNN. For the training and validation phases, we use data spanning from January 14, 2023, to December 16, 2023. Subsequently, data from December 17, 2023, to January 1, 2024, is exclusively reserved for testing and evaluating the models’ predictive performance.
Furthermore, from an engineering perspective, the training and implementation of neural network models are often constrained by limited computational resources, which are typically not as ample as those in computer science. Consequently, engineering applications necessitate relatively simple models capable of effectively handling challenges such as data noise and convexity. To ensure the selection of the most suitable model, it is essential to conduct comparisons under consistent settings (i.e., the same model complexity). This approach ensures fairness, as we evaluate not only accuracy but also other critical factors such as computational efficiency, which can be quantified by FLOPs. The comprehensive discussion of results is accessible in Appendix B.
IV-A1 Computational Efficiency Analysis
We leverage FLOPs as a measure to assess the computational efficiency of models. As depicted in Table I, it is evident that among the models compared, ICLRNN demonstrates the most efficient computational performance as a result of its lowest FLOPs. Consequently, this superior efficiency positions ICLRNN as the top-performing model within this context when considering computational resources.
Table I: Model complexity with layer size of (256, 256) for solar irradiance prediction.
Model Trainable Parameters FLOPs
RNN 199,682 399,362
LSTM 797,186 1,596,418
ICRNN 596,482 1,195,010
LRNN 330,754 2,498,562
ICLRNN (ours) 199,682 399,362
IV-A2 Non-Adversarial Robustness Analysis
In Fig. 2, we present results from the best state-of-the-art model, LRNN, alongside our model’s outcomes specifically for December 28, 2023. LRNN displays marked improvements compared to RNN and LSTM, yet it falls short in accurately predicting solar irradiance during sudden fluctuations. Additionally, during periods of minimal solar irradiance at the start and end of the day, LRNN exhibits slight fluctuations. Highlighted within Fig. 2, our ICLRNN outperforms LRNN, particularly in accurately predicting sudden changes in solar irradiance. This highlights the superior predictive capability of ICLRNN, showcasing its effectiveness in capturing and predicting nuanced fluctuations within the solar irradiance data. Notably, these observations persist across all trials and are not specific solely to December 28, 2023.
Figure 2: LRNN (top) and ICLRNN (bottom) performances on 2023-12-28 for solar irradiance prediction.
@@ IV-A2 Non-Adversarial Robustness Analysis
IV-B MPC for Continuous Stirred Tank Reactor
MPC is a real-time optimization-based control scheme, where the first element of the optimal input trajectory will be executed in the system over the next sampling period, and the MPC optimization problem will be resolved at the subsequent sampling period iteratively until convergence.
@@ IV-B MPC for Continuous Stirred Tank Reactor
Note that in general, solving NN-MPC is time-consuming, since it is a non-convex optimization problem due to the nonlinearity and nonconvexity of neural network models. However, if we integrate an ICNN into a convex MPC problem, the MPC problem will remain convex. In the next subsection, we will demonstrate the improvement of the computational efficiency of MPC using the proposed ICLRNN in Section III.
IV-B1 Application to a Chemical Process
@@
We implement LMPC to a nonisothermal and well-mixed continuous stirred tank reactor (CSTR) featuring an irreversible second-order exothermic reaction. This reaction facilitates the conversion of reactant A into product B. The CSTR incorporates a heating jacket responsible for either supplying or extracting heat at a rate $Q$. The dynamic model of the CSTR is characterized by the ensuing material and energy balance equations [29, 31]:
@@
where $C_{A}$ denotes the concentration of reactant A, $F$ denotes the volumetric flow rate, $V$ denotes the volume of the reactant, $C_{A0}$ denotes the inlet concentration of reactant A, $k_{0}$ denotes the pre-exponential constant, $E$ denotes the activation energy, $R$ denotes the ideal gas constant, $T$ denotes the temperature, $T_{0}$ denotes the inlet temperature, $\triangle H$ denotes the enthalpy of reaction, $\rho_{L}$ denotes the constant density of the reactant, $C_{p}$ denotes the heat capacity, and $Q$ denotes the heat input rate. The detailed constant values and system setup are discussed in [29, 31], and are omitted here.
@@
Moreover, $\triangle Q=Q-Q_{s}$ and $\triangle C_{A0}=C_{A0}-C_{A0_{s}}$ denote the manipulated inputs within this system, representing the alterations in the heat input rate and the inlet concentration of reactant A respectively. $x^{T}=[C_{A}-C_{As},T-T_{s}]$ denotes the system states, where $T-T_{s}$ and $C_{A}-C_{As}$ represent the deviations in the temperature of the reactor and the concentration of A from their respective steady-state values. $u^{T}=[\triangle C_{A0},\triangle Q]$ denotes the control actions. The primary objective of the controller involves operating the CSTR at $(C_{As},T_{s})$, which is the unstable equilibrium point, by manipulating $\triangle Q$ and $\triangle C_{A0}$ respectively. This manipulation is executed via the LMPC strategy outlined in Eq. (3) by ultimately stabilizing the state at its target steady-state.
@@
We conduct open-loop simulations for the CSTR of Eq. (4). These simulations aim to collect data that mimics the real-world data for training purposes (i.e., as it is difficult to obtain real-world chemical plant data). All models undergo training and evaluation using uniform configurations, incorporating the Categorical Cross-entropy Loss, the Adam Optimizer, and the ReLU activation function except for LRNN. The LMPC reaches the convergence state when $|x_{1}|=|C_{A}-C_{As}|<0.1\leavevmode\nobreak\ kmol/m^{3}$ and $|x_{2}|=|T-T_{s}|<3\leavevmode\nobreak\ K$ simultaneously. The initial conditions are chosen within the stability region, which is defined as $01060x_{1}^{2}+44x_{1}x_{2}+0.52x_{2}^{2}-372=$ 0 (i.e., an ellipse in state space). For a comprehensive evaluation of model performance across various hypothesis space sizes and initial conditions, detailed outcomes and discussions are presented in Appendix C.
IV-B2 Non-NN methods Performance
@@
In addition to NN-based methods, we investigate various data-driven models to provide a more comprehensive analysis, including non-NN methods such as the linear state-space model and the nonlinear autoregressive network with exogenous inputs (NARX), which have been widely used for modeling nonlinear dynamic systems for decades. As depicted in Table II, both the linear state-space model and the NARX model fail to outperform even a standard RNN.
Table II: Non-NN Methods Performance.
Model MSE
Linear State-space Model $1.17\times 10^{-3}$
NARX $6.37\times 10^{-4}$
Standard RNN $9.77\times 10^{-5}$
IV-B3 Model Computational Efficiency
To compare the computational efficiency in modeling the system dynamics in Eq. (4) across models, we quantify their FLOPs. Table III illustrates that ICLRNN has the lowest FLOPs, thereby outperforming or matching state-of-the-art methods. Similarly to the results shown in previous tasks, despite LRNN having a relatively modest count of trainable parameters, its high FLOPs indicate substantial computational demands, which is about 6 times more than ICLRNN. Conversely, ICRNN encounters challenges with exploding gradients, particularly when the network complexity exceeds a certain threshold.
Table III: Model complexity with layer size of (256, 256) for CSTR.
Model Trainable Parameters FLOPs
RNN 198,658 406,548
LSTM 793,090 1,597,460
ICRNN 596,482 1,204,233
LRNN 329,730 2,505,748
ICLRNN (ours) 198,658 406,548
IV-B4 Non-Adversarial Robustness Analysis
@@
To scrutinize the non-adversarial robustness of ICLRNN in modeling the system dynamics in Eq. (4), we assess the model’s sensitivity with respect to test accuracy when exposed to a series of perturbed inputs influenced by Gaussian noise. As illustrated in Fig. 3, the findings, averaging across 5 trials, reveal that ICLRNN consistently outperforms or matches the performance of state-of-the-art models across varying levels of noise.
Figure 3: Model performance w.r.t different degree of Gaussian noise for CSTR.
IV-B5 Computational Efficiency of ICLRNN-based LMPC
@@
Incorporating an ICNN into the aforementioned LMPC in Eq. (3) transforms it into a convex optimization problem. Subsequently, PyIpopt, as a Python version of IPOPT [32], is used to address the LMPC problem. The integration time step is set at $h_{c}=10^{-4}\leavevmode\nobreak\ hr$, while the sampling period remains $\triangle=0.005\leavevmode\nobreak\ hr$. The design of the Lyapunov function, denoted as $V(x)=x^{T}Px$, involves configuring the positive definite matrix $P$ as $\big{[}\begin{smallmatrix}1060&22\\ 22&0.52\end{smallmatrix}\big{]}$, which ensures the convexity of the LMPC.
We execute LMPC for a fixed timeframe for various models and repeat with various initial conditions to evaluate the performance of different NN-based LMPC. By examining Fig. 4, it is evident that the ICLRNN-based LMPC converges, whereas other NN-based LMPC models do not. The fastest convergence speed highlights the superior computational efficiency of the ICLRNN-based LMPC.
@@
Figure 4: LMPC converging path in a fixed timeframe with an initial condition at $[-1.5\leavevmode\nobreak\ kmol/m^{3},70\leavevmode\nobreak\ K]$.
V Limitations and Future Works
During our experiments, we observed that ICRNN faces challenges related to the exploding gradient problem, prompting us to discuss this issue and its relevance to the exploding and vanishing gradients phenomenon. Fortunately, as a beneficial side-effect of addressing our aforementioned objectives, our proposed method exhibits reduced susceptibility to the exploding gradient problem to a certain extent, attributable to its incorporation of Lipschitz constraints.
Subsequently, we further investigated the potential of ICLRNN by delving into the Long Range Arena benchmark [21], encompassing tasks involving long sequence modeling such as byte-level text sentiment classification and pixel sequence image classification.

### v4

@@ 3.2.2 MPC formulation
where $S(\triangle)$ denotes the set of piecewise constant functions with period $\triangle$, $\tilde{x}$ denotes the predicted state trajectory, and $N$ denotes the number of sampling periods in the prediction horizon. The objective function $\mathcal{L}$ outlined in Eq. (3a) integrates a cost function $J$ that depends on both the control actions $u$ and the system states $x$. The system dynamic function $F_{nn}(\tilde{x}(t),u(t))$ expressed in Eq. (3b) is parameterized by an RNN (e.g., the proposed ICLRNN in this work). Eq. (3c) encapsulates the constraint function $U$, delineating feasible control actions. Eq. (3d) establishes the initial condition $\tilde{x}(t_{k})$ in Eq. (3b), referring to the state measurement at $t=t_{k}$. Lastly, Eq. (3e) represents the Lyapunov-based constraint $V$, ensuring closed-loop stability within LMPC by mandating a decrease in the value of $V(x)$ (i.e., Lyapunov functions $V(x)$ are analogous to energy functions in some sense that the system is stable at the lowest energy point where $V(x)=0$ at $x=0$).
It is important to highlight that a convex neural network-based LMPC remains convex even when making multi-step ahead predictions, where the prediction horizon is greater than one [9]. This statement holds true if we incorporate an inherently input convex neural network (e.g., ICLRNN) to the LMPC, as described in Eq. (3b), and if we guarantee that the cost function $J$ in Eq. (3a) is input convex throughout the task.
4 Empirical evaluation
@@
Drawing from the above considerations, we meticulously choose two case studies to showcase the computational efficiency and non-adversarial robustness of ICLRNN. These include implementing an NN-MPC for a chemical reactor and predicting solar irradiance for Solar PV systems using real-time data from LHT Holdings plants in Singapore. It is worth noting that ablation studies are conducted to ascertain the optimal hyperparameter values for all subsequent tasks. Moreover, given that our research encompasses both Lipschitz-constrained NNs and Input Convex NNs, we meticulously select the most suitable approach from these categories and ensure consistency by employing a unified set of baseline methods across both tasks. Our primary aim, however, is to showcase the superiority of our proposed method over any alternatives within these two distinct families. Details on the computing resources are provided in Appendix D.
@@
4.1 Application to real-time optimization of continuous stirred tank reactor
@@
MPC is a real-time optimization-based control scheme, where the first element of the optimal input trajectory will be executed in the system over the next sampling period, and the MPC optimization problem will be resolved at the subsequent sampling period iteratively until convergence. Note that in general, solving NN-MPC is time-consuming, since it is a non-convex optimization problem due to the nonlinearity and nonconvexity of neural network models. However, if we integrate an ICNN into a convex MPC problem, the MPC problem will remain convex. In the next subsection, we demonstrate the improvement of the non-adversarial robustness in chemical process modeling and computational efficiency of NN-MPC using the proposed ICLRNN.
4.1.1 Nonlinear chemical process modeling
@@
We implement MPC to a nonisothermal and well-mixed continuous stirred tank reactor (CSTR) featuring an irreversible second-order exothermic reaction. This reaction facilitates the conversion of reactant A into product B. The CSTR incorporates a heating jacket responsible for either supplying or extracting heat at a rate $Q$. The dynamic model of the CSTR is characterized by the ensuing material and energy balance equations:
@@
where $C_{A}$ denotes the concentration of reactant A, $F$ denotes the volumetric flow rate, $V$ denotes the volume of the reactant, $C_{A0}$ denotes the inlet concentration of reactant A, $k_{0}$ denotes the pre-exponential constant, $E$ denotes the activation energy, $R$ denotes the ideal gas constant, $T$ denotes the temperature, $T_{0}$ denotes the inlet temperature, $\triangle H$ denotes the enthalpy of reaction, $\rho_{L}$ denotes the constant density of the reactant, $C_{p}$ denotes the heat capacity, and $Q$ denotes the heat input rate. The detailed parameter values are listed in Table 2 in Appendix B.
@@
Moreover, $x^{T}=[C_{A}-C_{As},T-T_{s}]$ denotes the system states, where $T-T_{s}$ and $C_{A}-C_{As}$ represent the deviations in the temperature of the reactor and the concentration of A from their respective steady-state values. $u^{T}=[\triangle C_{A0},\triangle Q]$ denotes the control actions, where $\triangle Q=Q-Q_{s}$ and $\triangle C_{A0}=C_{A0}-C_{A0_{s}}$ denote the manipulated inputs within this system, representing the alterations in the heat input rate and the inlet concentration of reactant A, respectively. The objective of the MPC of Eq. (3) involves operating the CSTR at $(C_{As},T_{s})$, which is the unstable equilibrium point, by manipulating $\triangle Q$ and $\triangle C_{A0}$ respectively. In summary, the inputs consist of $[T_{t}-T_{s},C_{A_{t}}-C_{As},Q_{t},C_{{A0}_{t}}]$ at the current time step $t$. The outputs entail the state trajectory $[T_{t+1,\ldots,n},C_{A_{t+1,\ldots,n}}]$ over the subsequent $n$ time steps (i.e., representing a one-to-many sequence problem).
@@
We conduct open-loop simulations for the CSTR of Eq. (4) using explicit Euler method. These simulations aim to collect data that mimics the real-world data for training purposes (i.e., as it is difficult to obtain real-world chemical plant data). All models undergo training and evaluation using uniform configurations, incorporating the mean-squared loss (MSE), the Adam Optimizer, and the ReLU activation function. The MPC reaches the convergence state when $|x_{1}|=|C_{A}-C_{As}|<0.1~{}kmol/m^{3}$ and $|x_{2}|=|T-T_{s}|<3~{}K$ simultaneously. The initial conditions are chosen within the stability region, which is defined as $01060x_{1}^{2}+44x_{1}x_{2}+0.52x_{2}^{2}-372=$ 0 (i.e., an ellipse in state space). For a comprehensive evaluation of model performance across various hypothesis space sizes and initial conditions, detailed outcomes and discussions are presented in Appendix B.
4.1.2 Non-NN methods performance
@@
In addition to NN-based methods, we investigate various data-driven models to provide a more comprehensive analysis, including non-NN methods such as the linear state-space model and the nonlinear autoregressive network with exogenous inputs (NARX), which have been widely used for modeling nonlinear dynamic systems for decades. Both the linear state-space model (i.e., testing MSE of $1.17\times 10^{-3}$) and the NARX model (i.e., testing MSE of $6.37\times 10^{-4}$) fail to outperform even a standard RNN (i.e., testing MSE of $9.77\times 10^{-5}$).
4.1.3 Performance analysis for CSTR modeling
Based on our experiments, we observed that a testing MSE on the order of $10^{-3}$ is indicative of satisfactory performance, rendering the model suitable for applications such as NN-MPC (i.e., NN-MPC can converge). To compare the computational efficiency in modeling the system of Eq. (4), we quantify their Floating Point Operations per Second (FLOPs). Table 1 illustrates that ICLRNN has the lowest FLOPs, thereby outperforming or matching state-of-the-art methods. It is crucial to emphasize that the FLOPs presented in Table 1 are calculated for models that share identical architectures, including the same number of layers and neurons.
Table 1: Model complexity in terms of FLOPs.
Model FLOPs (CSTR) FLOPs (Solar PV)
RNN 406,548 399,362
LSTM 1,597,460 1,596,418
ICRNN 1,204,233 1,195,010
LRNN 2,505,748 2,498,562
ICLRNN (ours) 406,548 399,362
@@ 4.1.3 Performance analysis for CSTR modeling
To scrutinize the non-adversarial robustness of ICLRNN in modeling the system dynamics in Eq. (4), we assess the model’s sensitivity with respect to test accuracy when exposed to a series of perturbed inputs influenced by Gaussian noise. As illustrated in Fig. 1(a), the findings, averaging across 5 trials, reveal that ICLRNN consistently outperforms or matches the performance of state-of-the-art models across varying levels of noise. These findings indicate ICLRNN’s non-adversarial robustness across diverse hypothesis space sizes with respect to different Gaussian noise levels. Despite LRNN having a relatively modest testing MSE against different Gaussian noise levels, its high FLOPs indicate substantial computational demands, which are about 6 times greater than ICLRNN. This computational intensity stands as a significant drawback despite its competitive performance. Notably, ICLRNN exhibits superior computational efficiency, as evidenced by its notably lower number of FLOPs compared to other models. Moreover, ICRNN encounters challenges when the hypothesis space expands, succumbing to the issue of the exploding gradient problem, which makes it difficult to train.
4.1.4 Performance analysis for CSTR-based LMPC
@@
Next, incorporating an ICNN into the aforementioned LMPC in Eq. (3) transforms it into a convex optimization problem. Subsequently, PyIpopt, as a Python version of IPOPT [22], is used to address the LMPC problem. The integration time step is set at $h_{c}=1\times 10^{-4}~{}hr$, while the sampling period remains $\triangle=5\times 10^{-3}~{}hr$. The design of the Lyapunov function, denoted as $V(x)=x^{T}Px$, involves configuring the positive definite matrix $P$ as $\big{[}\begin{smallmatrix}1060&22\\ 22&0.52\end{smallmatrix}\big{]}$, which ensures the convexity of the LMPC. We execute LMPC for a fixed timeframe for various models and repeat with various initial conditions to evaluate the performance of different NN-based LMPC. Fig. 1(b) shows one of the initial conditions, from which it is evident that the ICLRNN-based LMPC converges, whereas other NN-based LMPC models do not. The fastest convergence speed highlights the superior computational efficiency of the ICLRNN-based LMPC. Moreover, it is important to highlight that ICLRNN-based LMPC consistently demonstrates the fastest convergence speed across various initial conditions, thereby surpassing the state-of-the-art methodologies. Performance with respect to other initial conditions are discussed in Appendix B.
(a) Model performance w.r.t different degree of Gaussian noise for CSTR.
@@
(b) LMPC converging path in a fixed timeframe with an initial condition at $[-1.5~{}kmol/m^{3},70~{}K]$.
Figure 1: Performance of CSTR-based LMPC.
4.2 Application to real-world solar irradiance prediction
In this task, we aim to forecast solar irradiance using real-time data for LHT Holdings, a wood pallet manufacturing industry based in Singapore. Before the installation of the Solar photovoltaic (PV) system, LHT Holdings relied solely on the main utility grid to fulfill its energy needs. The Solar PV system was successfully installed by 10 Degree Solar in late 2022 (i.e., a detailed manufacturing pipeline of LHT Holdings and the solar PV system are available in Appendix C). This transition to solar energy was primarily driven by the aim of minimizing manufacturing costs. Three significant uncertainties were identified within the industry, namely solar irradiance affecting the Solar PV system’s efficiency, the dynamic real-time pricing of energy sourced from the main utility grid, and the fluctuating energy demand stemming from unexpected customer orders.
Currently, the industrial facility primarily relies on Solar PV as its main energy source, with the main utility grid serving as a backup. Excess solar energy is stored for later use. Accurate solar irradiance forecasting is crucial for efficient energy planning and cost reduction. Using sensors installed by the Solar Energy Research Institute of Singapore (SERIS) and Singapore Institute of Manufacturing Technology (SIMTech), including irradiance, humidity, wind, and temperature sensors, alongside data similar to that used by [23], we predict minute-by-minute solar irradiance to optimize energy dispatch. In this time-series forecasting task, all models undergo training and evaluation using identical settings. These settings encompass MSE as the loss function, the Adam Optimizer, the ReLU activation function. For the training and validation phases, we use data spanning from January 14, 2023, to December 16, 2023. Subsequently, data from December 17, 2023, to January 1, 2024, is exclusively reserved for testing and evaluating the models’ predictive performance.
Furthermore, from an engineering perspective, the training and implementation of neural network models are often constrained by limited computational resources, which are typically not as ample as those in computer science. Consequently, engineering applications necessitate relatively simple models capable of effectively handling challenges such as data noise and convexity. To ensure the selection of the most suitable model, it is essential to conduct comparisons under consistent settings (i.e., the same model complexity). This approach ensures fairness, as we evaluate not only accuracy but also other critical factors such as computational efficiency, which can be quantified by FLOPs. The comprehensive discussion of results is accessible in Appendix C.
4.2.1 Performance analysis for solar irradiance forecasting
We leverage FLOPs as a measure to assess the computational efficiency of models. As depicted in Table 1, it is evident that among all the models, ICLRNN demonstrates the most efficient computational performance as a result of its lowest FLOPs. Consequently, this superior efficiency positions ICLRNN as the top-performing model when considering computational resources.
In Fig. 2(a), we present results from the best state-of-the-art model, LRNN, alongside our model’s outcomes specifically for December 28, 2023. The LRNN, constrained by Lipschitz continuity, exhibits significantly improved performance compared to the RNN and LSTM models. However, it still encounters challenges in accurately predicting solar irradiance during sudden fluctuations, and its excessively high FLOPs signify its computational intensity (i.e., 2,498,562 for LRNN compared to 399,362 for ICLRNN). Moreover, slight fluctuations in the LRNN predictions occur during periods of minimal solar irradiance at the start and end of the day. Highlighted within Fig. 2(b), our ICLRNN outperforms LRNN and other state-of-the-art models, particularly in accurately predicting sudden changes in solar irradiance, which highlights the superior predictive capability of ICLRNN, showcasing its effectiveness in capturing and predicting nuanced fluctuations within the solar irradiance data. Notably, these observations persist across all trials and are not specific solely to December 28, 2023.
(a) LRNN performances on 2023-12-28 for solar irradiance prediction.
(b) ICLRNN performances on 2023-12-28 for solar irradiance prediction.
Figure 2: Performance of solar irradiance forecasting.
5 More applications and discussions